Jack is twice as old as Jill. In 10 years, Jill's age will be 3/4 of jack's age. How old are they now.

Found 2 solutions by vleith, checkley77:
Answer by vleith(2983)

Jack = 2* Jill
Jack - 2Jill = 0
(3/4)(Jack+10) = Jill+10
Jack+10 = (4/3)Jill + 40/3
Jack = (4/3)Jill +10/3
3Jack = 4Jill + 10
3Jack - 4Jill = 10
Then solve using substitution or elimination. See here --> http://www.1728.com/unknwn2.htm
Jack 10, Jill 5

Answer by checkley77(12844)

3/4(2x+10)=x+10
6x/4+30/4=x+10
(6x+30)/4=x+10
6x+30=4x+40
6x-4x=40-30
2x=10
x=10/2
x=5 Jill's age now.
2*5=10 Jack's age now.
proof
3/4(10+10)=5+10
3/4*20=15
60/4=15
15=15